The length of a rectangle is 8 cm mor than 3 times its width. the perimeter of the rectangle is 64 cm.
B. What is the area of the rectangle?

Answer by checkley71(8403)

LENGTH=3W+8
PERIMETER=2L+2W
64=2(3W+8)+2W
64=6W+16+2W
8W=64-16
8W=48
W=48/8
W=6 ANSWER FOR THE WIDTH.
L=3*6+8
L=18+8
L=26 ANSWER FOR THE LENGTH.
PROOF
2*6+2*26=64
12+52=64
64=64
---------------------------
AREA=L*W
AREA=26*6
AREA=156 CM^2